Rotational mechanics concept torque formula in the view of dimensional formula

Grade:9

1 Answers

Vikas TU
7 years ago
Torque is given by the formulae:
T =  F x r 
where r is the perpendicular distance from the axis of rotation.
F = ma = mv/t = mr/t^2
T = mr^2/t^2
Dimension would be:
ML^2T^-2.